OpenStudy (anonymous):

I could use some help to factor this question. The area of a painting is given by the trinomial x²+4x-21. What are the possible demensions of the painting?

OpenStudy (anonymous):

(x-3)(x-7) is that trinomial factored.

OpenStudy (anonymous):

(x+7)*

OpenStudy (anonymous):

THanks! Are you sure?

jhonyy9 (jhonyy9):

(x-3)(x+7)=x^2 -3x +7x -21 = x^2 +4x -21 so is right sure